OpenStudy (anonymous):

Solve by factoring x^2-9x=-20

OpenStudy (anonymous):

x^2 - 9x + 20 = 0 (x-5)(x-4)=0

OpenStudy (anonymous):

x = 5, x = 4

OpenStudy (anonymous):

Factor it: /[x^2-9x+20/] /[(x-4)(x-5)/]